OpenStudy (anonymous):

The sum of two numbers is 60. Their difference is 12. What is the smaller number?

OpenStudy (eyust707):

okay so with this information we can write down some expressions.

OpenStudy (eyust707):

"the sum of two numbers is 60"

OpenStudy (eyust707):

"sum" means addition

OpenStudy (eyust707):

so that sentence means: Some number plus another number equals 60.

OpenStudy (eyust707):

"some number" + "another number" = 60 in math, to simplify things we can represent "some number" and "another number" with letters. I like to use x and y: x + y =60

OpenStudy (eyust707):

"Their difference is 12." Difference usually refers to subtraction. So that sentence means "the difference of some number and another number is 12" or x - y =12

OpenStudy (eyust707):

so now we have: x - y = 12 x + y = 60

OpenStudy (eyust707):

do you know of a way to figure out what x or y is?

OpenStudy (anonymous):

No I dont

OpenStudy (eyust707):

we can use a method called substitution.

OpenStudy (eyust707):

lets start with the top equation

OpenStudy (eyust707):

x - y = 12

OpenStudy (eyust707):

if you add y to both sides what do you get?

OpenStudy (anonymous):

24

OpenStudy (eyust707):

x - y = 12 x = 12 + y

OpenStudy (eyust707):

do you see how i did that?

OpenStudy (anonymous):

oh yes

OpenStudy (eyust707):

okay so now we know for sure that x equals "12 + y"

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

well that means that we can go back to our other equation and plug "12 + y" in for "x"

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

x = "12 + y" and from before, x + y = 60 that means that "12 + y" + y = 60

OpenStudy (eyust707):

(usually in math we dont write quotes, instead we use parenthesis) (12 + y) + y = 60 12 + y + y =60

OpenStudy (eyust707):

if i have y + y how many y's do I have?

OpenStudy (anonymous):

24

OpenStudy (eyust707):

y + y = ??

OpenStudy (anonymous):

12

OpenStudy (eyust707):

y + y = 2y

OpenStudy (eyust707):

one y plus another y is two y's

OpenStudy (eyust707):

=P

OpenStudy (eyust707):

12 + y + y =60 12 + 2y = 60

OpenStudy (anonymous):

So Wat Would The Smaller number be?

OpenStudy (eyust707):

well we gotta figure out x and y to know which one is smaller useing this you can figure out y! 12 + 2y = 60

OpenStudy (eyust707):

minus 12 from both sides

OpenStudy (eyust707):

then divide both sides by 2

OpenStudy (eyust707):

what do you get for y?

OpenStudy (eyust707):

y = 24 x + y = 60 x + 24 = 60 x = 36

OpenStudy (eyust707):

the smaller one is 24